# Bonsai: Gradient-free Graph Distillation for Node Classification

### v1

@@ 3.4 Properties of Bonsai
Complexity analysis: The computation compelxity of Bonsai is $\mathcal{O}(\lvert\mathcal{V}\rvert+\lvert\mathcal{E}\rvert)$. The detailed derivation is provided in App. A.4.
CPU-bound and Parallelizable: Bonsai does not involve learning any parameters through gradient descent or its variants. Hence, the entire process is CPU-bound. Furthermore, all steps are embarrassingly parallelizable leading distillation of datasets with hundreds of millions of edges within minutes.
Independence from model-architecture and hyper-parameters: Unlike majority of existing distillation algorithms that require knowledge of the Gnn architecture and all training hyper-parameters, Bonsai only requires approximate knowledge of the number of layers $L$. An approximate knowledge is sufficient since the influence of far-away neighbors gets squashed (Topping et al., 2022). For the same reason, Gnns are typically run with $L\leq 3$. Hence, we recommend setting $L\sim 2$.
@@ 4 Experiments
  • Enhanced Efficiency: On average, Bonsai is 350, 22, and 1500 times faster than the state-of-the-art baselines GCond, Gdem and Gcsr, respectively, in distilling datasets.
  • Increased Robustness: Unlike existing methods that require tuning distillation-specific hyper-parameters for each combination of Gnn architecture, dataset, and compression ratio, our approach surpasses all baselines using a single set of parameters across all scenarios.
Our implementation is available at https://anonymous.4open.science/r/bonsai.
@@ 4.1 Experimental Setup
The specifics of our experimental setup, including hardware and software environment, and hyper-parameters are detailed in App. B. For the baseline algorithms, we use the code shared by their respective authors. We conduct each experiment $5$ times and report the means and standard deviations. Across all datasets, we maintain a train-validation-test split ratio of $60:20:20$.
@@
Datasets: Table 3 lists the benchmark datasets used. The same datasets were also used in GCond, Gdem and Gcsr.
@@
Baselines: Table 1 lists all the graph distillation algorithms for node classification. We compare Bonsai with GCond (Jin et al., 2021), Gdem (Liu et al., 2024a), and Gcsr (Liu et al., 2024b). We omit Sfgc (Zheng et al., 2023) and Sgdd (Yang et al., 2023) since both Gdem and Gcsr have been shown to outperform them. Among non-neural, baselines, we compare with selecting the induced subgraph spanned by a random selection of nodes, and Herding (Welling, 2009).
@@
Metrics: Prediction quality is measured through Accuracy on the test set, i.e., the percentage of correct predictions. Compression ratio is quantified as $S_{r}=\frac{\text{size of distilled dataset in bytes}}{\text{size of full dataset in bytes}}$.
@@
Table 4: Accuracy achieved by the various baselines on benchmark datasets across various compression ratios over byte consumption (denoted as $\mathbf{S_{r}(\%)}$), , on the Gcn architecture. The best and the second-best accuracies in each row are highlighted by dark and lighter shades of Green, respectively. OOT indicates the scenario where the algorithm failed to distill within 5 hours.
@@
Dataset $\mathbf{S_{r}(\%)}$ Random Herding GCond GDEM GCSR Bonsai Full
0.5 $39.90{\scriptstyle\pm 1.46}$ $45.61{\scriptstyle\pm 0.01}$ $77.30{\scriptstyle\pm 0.30}$ $57.49{\scriptstyle\pm 6.87}$ $74.83{\scriptstyle\pm 0.75}$ $83.95{\scriptstyle\pm 0.39}$
Cora 1 $27.75{\scriptstyle\pm 2.05}$ $52.07{\scriptstyle\pm 0.00}$ $77.30{\scriptstyle\pm 0.31}$ $69.32{\scriptstyle\pm 4.71}$ $77.56{\scriptstyle\pm 0.74}$ $85.76{\scriptstyle\pm 0.24}$ $88.56{\scriptstyle\pm 0.18}$
3 $52.26{\scriptstyle\pm 1.69}$ $67.60{\scriptstyle\pm 0.00}$ $81.73{\scriptstyle\pm 0.48}$ $81.70{\scriptstyle\pm 3.10}$ $77.20{\scriptstyle\pm 0.48}$ $86.38{\scriptstyle\pm 0.22}$
0.5 $33.90{\scriptstyle\pm 2.16}$ $22.82{\scriptstyle\pm 0.00}$ $74.17{\scriptstyle\pm 0.68}$ $70.05{\scriptstyle\pm 2.40}$ $67.03{\scriptstyle\pm 0.61}$ $77.00{\scriptstyle\pm 0.15}$
CiteSeer 1 $44.90{\scriptstyle\pm 2.32}$ $49.10{\scriptstyle\pm 0.02}$ $77.62{\scriptstyle\pm 0.71}$ $72.48{\scriptstyle\pm 2.13}$ $74.77{\scriptstyle\pm 0.78}$ $77.03{\scriptstyle\pm 0.33}$ $78.53{\scriptstyle\pm 0.15}$
3 $44.50{\scriptstyle\pm 1.27}$ $67.69{\scriptstyle\pm 0.01}$ $77.02{\scriptstyle\pm 0.22}$ $76.20{\scriptstyle\pm 0.55}$ $77.27{\scriptstyle\pm 0.28}$ $75.89{\scriptstyle\pm 0.26}$
0.5 $62.58{\scriptstyle\pm 0.25}$ $78.29{\scriptstyle\pm 0.00}$ $80.63{\scriptstyle\pm 1.20}$ $80.72{\scriptstyle\pm 0.92}$ $79.43{\scriptstyle\pm 0.25}$ $87.27{\scriptstyle\pm 0.03}$
Pubmed 1 $79.19{\scriptstyle\pm 0.09}$ $78.59{\scriptstyle\pm 0.00}$ $79.92{\scriptstyle\pm 0.00}$ $80.80{\scriptstyle\pm 1.07}$ $79.11{\scriptstyle\pm 0.15}$ $87.08{\scriptstyle\pm 0.04}$ $87.22{\scriptstyle\pm 0.00}$
3 $82.50{\scriptstyle\pm 0.09}$ $78.09{\scriptstyle\pm 0.00}$ $77.00{\scriptstyle\pm 0.15}$ $81.07{\scriptstyle\pm 0.90}$ $79.94{\scriptstyle\pm 0.16}$ $87.64{\scriptstyle\pm 0.09}$
0.5 $44.78{\scriptstyle\pm 0.00}$ $47.98{\scriptstyle\pm 0.01}$ $44.06{\scriptstyle\pm 1.05}$ $46.25{\scriptstyle\pm 1.02}$ $46.41{\scriptstyle\pm 0.00}$ $48.73{\scriptstyle\pm 0.27}$
@@
3 $46.56{\scriptstyle\pm 0.01}$ $46.54{\scriptstyle\pm 0.01}$ $46.04{\scriptstyle\pm 1.88}$ $47.35{\scriptstyle\pm 0.98}$ OOT $49.66{\scriptstyle\pm 0.27}$
0.5 $42.01{\scriptstyle\pm 0.01}$ $53.37{\scriptstyle\pm 0.00}$ $52.63{\scriptstyle\pm 0.63}$ $54.73{\scriptstyle\pm 0.66}$ OOT $58.49{\scriptstyle\pm 0.17}$
Ogbn-arxiv 1 $49.27{\scriptstyle\pm 0.64}$ $54.91{\scriptstyle\pm 0.00}$ $53.49{\scriptstyle\pm 0.63}$ $51.45{\scriptstyle\pm 1.14}$ OOT $58.35{\scriptstyle\pm 0.09}$ $68.97{\scriptstyle\pm 0.10}$
3 $51.11{\scriptstyle\pm 0.19}$ $57.28{\scriptstyle\pm 0.00}$ $53.01{\scriptstyle\pm 0.64}$ $53.37{\scriptstyle\pm 1.04}$ OOT $64.31{\scriptstyle\pm 0.06}$
0.5 $36.00{\scriptstyle\pm 4.09}$ OOT $38.94{\scriptstyle\pm 0.79}$ $90.51{\scriptstyle\pm 0.55}$ OOT $80.33{\scriptstyle\pm 0.46}$
Reddit 1 $38.55{\scriptstyle\pm 2.00}$ OOT $43.98{\scriptstyle\pm 0.35}$ $90.63{\scriptstyle\pm 0.84}$ OOT $85.65{\scriptstyle\pm 0.08}$ $92.14{\scriptstyle\pm 0.04}$
3 $44.97{\scriptstyle\pm 2.97}$ OOT $48.78{\scriptstyle\pm 0.83}$ $85.75{\scriptstyle\pm 0.80}$ OOT $88.90{\scriptstyle\pm 0.07}$
@@ 4.2 Prediction Accuracy
Table 4 presents the accuracies obtained by Bonsai and the baselines on Gcn. Several insights emerge from this experiment. Firstly, Bonsai achieves the best accuracy in 14 out of 18 scenarios, with substantial improvements ($\geq 5\%$) over the second-best performer in several cases. This demonstrates that the unsupervised gradient-agnostic approach adopted by Bonsai does not come at the cost of accuracy. Gdem emerges as the second best perfomer. Gcsr fails to complete in the two largest datasets of Ogbn-arxiv and Reddit since it performs training on the full dataset $100$ times to distill, which exceeds 5 hours.
Second, even at extreme compression ratio of $0.5\%$, it is possible to achieve reasonable accuracy levels. We highlight that past studies have denoted compression ratio in terms of number of nodes (recollect Table 2), which is misleading, since they generate fully connected weighted graphs, and thus real-compression in terms of bytes consumption, is lower.
@@ 4.3 Cross-architecture Generalization
Gdem and Bonsai are both model-agnostic, meaning their distilled datasets are independent of the downstream Gnn architecture used. In contrast, GCond and Gcsr produce architecture-specific distillation datasets, requiring separate training for each architecture. This distinction raises two important questions: First, how well do the distilled datasets generated by Gdem and Bonsai generalize across different architectures? Second, if we apply the distillation datasets produced by GCond and Gcsr for Gcn to other architectures, how significantly does it impact performance? We explore these questions in Table 5.
@@
Consistent with the earlier trend, Bonsai continues to outperform all baselines. We further observe that the performance gap between Bonsai and the baselines is wider in Gat and Gin compared to Gcn (Table 4). For GCond and Gcsr, this wider gap is not surprising since they are trained on the gradients of Gcn, which are expected to differ from those of Gin and Gat. Gin uses a Sum-Pool aggregation, while Gat employs attention to effectively dampen unimportant edges, with a potentially different eigen spectrum - the property that Gdem attempts to preserve. The approach of Bonsai, on the other hand, aligns more directly with the computational structure of Gnns. It focuses on analyzing the input space of computation trees and aims to represent as much of this space as possible within the given budget. This approach of retaining essential information, rather than mimicking a pre-determined proxy function assumed to be important for Gnns, leads to superior results.
Table 5: Accuracies achieved by the various baselines on Gat and Gin at $1\%$ compression.

### v3

@@ 3.3 Properties of Bonsai
Complexity analysis: The computation complexity of Bonsai is $\mathcal{O}(\lvert\mathcal{V}\rvert+\lvert\mathcal{E}\rvert)$ (details in App. A.5).
CPU-bound and Parallelizable: Bonsai does not involve learning any parameters through gradient descent or its variants. Hence, the entire process is CPU-bound. Furthermore, all steps are embarrassingly parallelizable leading condensation of datasets with hundreds of millions of edges within minutes.
Independence from model-architecture and hyper-parameters: Unlike majority of existing condensation algorithms that require knowledge of the Gnn architecture and all training hyper-parameters, Bonsai only requires approximate knowledge of the number of layers $L$.
@@ 4 Experiments
  • Enhanced computation and energy efficiency: On average, Bonsai is at least 7 times faster and 17 times more energy efficient than the state-of-the-art baselines.
  • Increased robustness: Unlike existing methods that require tuning condensation-specific hyper-parameters for each combination of Gnn architecture, dataset, and compression ratio, Bonsai achieves superior performance using a single set of parameters across all scenarios.
Our implementation is available at https://github.com/idea-iitd/Bonsai.
@@ 4.1 Experimental Setup
MAG240M (Hu et al., 2021) 1,398,159 26,434,726 153 768
The specifics of our experimental setup, including hardware and software environment, and hyper-parameters are detailed in App. B. For the baseline algorithms, we use the code shared by their respective authors. We conduct each experiment $5$ times and report the means and standard deviations.
Datasets: Table 3 lists the benchmark datasets used.
@@
Baselines: Table 1 lists all the graph condensation algorithms for node classification. We omit Sfgc (Zheng et al., 2023) and Sgdd (Yang et al., 2023) as baselines since both Gdem and Gcsr have been shown to outperform them. Among non-neural, baselines, we compare with selecting the induced subgraph spanned by a random selection of nodes, and Herding (Welling, 2009).
@@
Metrics: Prediction quality is measured through Accuracy on the test set, i.e., the percentage of correct predictions. Compression ratio is quantified as $S_{r}=\frac{\text{size of condensed dataset in bytes}}{\text{size of full dataset in bytes}}$.
@@
Table 4: Accuracy achieved by the various baselines on benchmark datasets across various compression ratios over byte consumption (denoted as $\mathbf{S_{r}(\%)}$), on the Gcn architecture. The best and the second-best accuracies in each row are highlighted by dark and lighter shades of Green, respectively. OOT indicates the scenario where the algorithm failed to condense within 5 hours.
@@
Dataset $\mathbf{S_{r}(\%)}$ Random Herding GCond GDEM GCSR EXGC GCSNTK GEOM Bonsai Full
0.5 $39.90{\scriptstyle\pm 1.46}$ $45.61{\scriptstyle\pm 0.01}$ $77.30{\scriptstyle\pm 0.30}$ $57.49{\scriptstyle\pm 6.87}$ $74.83{\scriptstyle\pm 0.75}$ $34.87{\scriptstyle\pm 0.83}$ $77.85{\scriptstyle\pm 0.76}$ $33.76{\scriptstyle\pm 0.96}$ $83.95{\scriptstyle\pm 0.39}$
Cora 1 $27.75{\scriptstyle\pm 2.05}$ $52.07{\scriptstyle\pm 0.00}$ $77.30{\scriptstyle\pm 0.31}$ $69.32{\scriptstyle\pm 4.71}$ $77.56{\scriptstyle\pm 0.74}$ $35.24{\scriptstyle\pm 0.60}$ $67.52{\scriptstyle\pm 0.77}$ $33.02{\scriptstyle\pm 0.92}$ $85.76{\scriptstyle\pm 0.24}$ $88.56{\scriptstyle\pm 0.18}$
3 $52.26{\scriptstyle\pm 1.69}$ $67.60{\scriptstyle\pm 0.00}$ $81.73{\scriptstyle\pm 0.48}$ $81.70{\scriptstyle\pm 3.10}$ $77.20{\scriptstyle\pm 0.48}$ $35.42{\scriptstyle\pm 0.77}$ $75.64{\scriptstyle\pm 0.82}$ $54.80{\scriptstyle\pm 1.89}$ $86.38{\scriptstyle\pm 0.22}$
0.5 $33.90{\scriptstyle\pm 2.16}$ $22.82{\scriptstyle\pm 0.00}$ $74.17{\scriptstyle\pm 0.68}$ $70.05{\scriptstyle\pm 2.40}$ $67.03{\scriptstyle\pm 0.61}$ $23.42{\scriptstyle\pm 0.98}$ $69.82{\scriptstyle\pm 0.68}$ $24.92{\scriptstyle\pm 0.89}$ $77.00{\scriptstyle\pm 0.15}$
CiteSeer 1 $44.90{\scriptstyle\pm 2.32}$ $49.10{\scriptstyle\pm 0.02}$ $77.62{\scriptstyle\pm 0.71}$ $72.48{\scriptstyle\pm 2.13}$ $74.77{\scriptstyle\pm 0.78}$ $23.67{\scriptstyle\pm 1.00}$ $69.22{\scriptstyle\pm 0.71}$ $28.03{\scriptstyle\pm 0.81}$ $77.03{\scriptstyle\pm 0.33}$ $78.53{\scriptstyle\pm 0.15}$
3 $44.50{\scriptstyle\pm 1.27}$ $67.69{\scriptstyle\pm 0.01}$ $77.02{\scriptstyle\pm 0.22}$ $76.20{\scriptstyle\pm 0.55}$ $77.27{\scriptstyle\pm 0.28}$ $25.07{\scriptstyle\pm 0.92}$ $64.26{\scriptstyle\pm 0.53}$ $33.48{\scriptstyle\pm 0.83}$ $75.89{\scriptstyle\pm 0.26}$
0.5 $62.58{\scriptstyle\pm 0.25}$ $78.29{\scriptstyle\pm 0.00}$ $80.63{\scriptstyle\pm 1.20}$ $80.72{\scriptstyle\pm 0.92}$ $79.43{\scriptstyle\pm 0.25}$ $45.77{\scriptstyle\pm 0.73}$ $53.04{\scriptstyle\pm 1.99}$ OOT $87.27{\scriptstyle\pm 0.03}$
Pubmed 1 $79.19{\scriptstyle\pm 0.09}$ $78.59{\scriptstyle\pm 0.00}$ $79.92{\scriptstyle\pm 0.00}$ $80.80{\scriptstyle\pm 1.07}$ $79.11{\scriptstyle\pm 0.15}$ $46.24{\scriptstyle\pm 0.43}$ $62.81{\scriptstyle\pm 1.32}$ OOT $87.08{\scriptstyle\pm 0.04}$ $87.22{\scriptstyle\pm 0.00}$
3 $82.50{\scriptstyle\pm 0.09}$ $78.09{\scriptstyle\pm 0.00}$ $77.00{\scriptstyle\pm 0.15}$ $81.07{\scriptstyle\pm 0.90}$ $79.94{\scriptstyle\pm 0.16}$ $47.62{\scriptstyle\pm 0.67}$ $67.72{\scriptstyle\pm 2.01}$ OOT $87.64{\scriptstyle\pm 0.09}$
0.5 $44.78{\scriptstyle\pm 0.00}$ $47.98{\scriptstyle\pm 0.01}$ $44.06{\scriptstyle\pm 1.05}$ $46.25{\scriptstyle\pm 1.02}$ $46.41{\scriptstyle\pm 0.00}$ $45.47{\scriptstyle\pm 0.85}$ $31.49{\scriptstyle\pm 0.75}$ OOT $48.73{\scriptstyle\pm 0.27}$
@@
3 $46.56{\scriptstyle\pm 0.01}$ $46.54{\scriptstyle\pm 0.01}$ $46.04{\scriptstyle\pm 1.88}$ $47.35{\scriptstyle\pm 0.98}$ OOT $48.44{\scriptstyle\pm 0.65}$ $36.58{\scriptstyle\pm 0.82}$ OOT $49.66{\scriptstyle\pm 0.27}$
0.5 $42.01{\scriptstyle\pm 0.01}$ $53.37{\scriptstyle\pm 0.00}$ $52.63{\scriptstyle\pm 0.63}$ $54.73{\scriptstyle\pm 0.66}$ OOT $60.66{\scriptstyle\pm 1.66}$ $61.55{\scriptstyle\pm 1.17}$ OOT $58.49{\scriptstyle\pm 0.17}$
Ogbn-arxiv 1 $49.27{\scriptstyle\pm 0.64}$ $54.91{\scriptstyle\pm 0.00}$ $53.49{\scriptstyle\pm 0.63}$ $51.45{\scriptstyle\pm 1.14}$ OOT $61.73{\scriptstyle\pm 1.43}$ $62.31{\scriptstyle\pm 0.88}$ OOT $58.35{\scriptstyle\pm 0.09}$ $68.97{\scriptstyle\pm 0.10}$
3 $51.11{\scriptstyle\pm 0.19}$ $57.28{\scriptstyle\pm 0.00}$ $53.01{\scriptstyle\pm 0.64}$ $53.37{\scriptstyle\pm 1.04}$ OOT $62.96{\scriptstyle\pm 1.33}$ $56.38{\scriptstyle\pm 0.79}$ OOT $64.31{\scriptstyle\pm 0.06}$
0.5 $36.00{\scriptstyle\pm 4.09}$ ${81.72}{\scriptstyle\pm{0.63}}$ $38.94{\scriptstyle\pm 0.79}$ $90.51{\scriptstyle\pm 0.55}$ OOT $78.47{\scriptstyle\pm 0.52}$ $37.15{\scriptstyle\pm 1.51}$ OOT $80.33{\scriptstyle\pm 0.46}$
Reddit 1 $38.55{\scriptstyle\pm 2.00}$ ${83.48}{\scriptstyle\pm{0.83}}$ $43.98{\scriptstyle\pm 0.35}$ $90.63{\scriptstyle\pm 0.84}$ OOT $81.69{\scriptstyle\pm 1.12}$ $38.87{\scriptstyle\pm 2.00}$ OOT $85.65{\scriptstyle\pm 0.08}$ $92.14{\scriptstyle\pm 0.04}$
3 $44.97{\scriptstyle\pm 2.97}$ ${88.51}{\scriptstyle\pm{0.13}}$ $48.78{\scriptstyle\pm 0.83}$ $85.75{\scriptstyle\pm 0.80}$ OOT OOM $47.48{\scriptstyle\pm 1.98}$ OOT $88.90{\scriptstyle\pm 0.07}$
0.5 $34.43{\scriptstyle\pm 0.11}$ $36.18{\scriptstyle\pm 0.10}$ OOT OOM OOT OOM OOT OOT $52.33{\scriptstyle\pm 0.05}$
MAG240M 1 $37.93{\scriptstyle\pm 0.08}$ $37.29{\scriptstyle\pm 0.06}$ OOT OOM OOT OOM OOT OOT $52.58{\scriptstyle\pm 0.33}$ $69.95{\scriptstyle\pm 3.52}$
3 $42.96{\scriptstyle\pm 0.43}$ $37.98{\scriptstyle\pm 0.09}$ OOT OOM OOT OOM OOT OOT $53.39{\scriptstyle\pm 0.07}$
@@ 4.2 Prediction Accuracy
Table 4 presents the accuracies obtained by Bonsai and the baselines on Gcn. Bonsai achieves the best accuracy in 12 out of 18 scenarios, with substantial improvements ($\geq 5\%$) over the second-best performer in several cases. This demonstrates that the unsupervised gradient-agnostic approach adopted by Bonsai does not come at the cost of accuracy. Gdem emerges as the second best perfomer. Gcsr fails to complete in the two largest datasets of Ogbn-arxiv and Reddit, since it trains on the entire dataset $100$ times to condense, which exceeds 5 hours. Geom is even slower since it trains on full train set $200$ times. All baselines except Random and Herding failed to finish within 5 hours on MAG240M.
@@ 4.3 Cross-architecture Generalization
Gdem and Bonsai are both model-agnostic, meaning their condensed datasets are independent of the downstream Gnn architecture used. In contrast, GCond and Gcsr produce architecture-specific condensation datasets, requiring separate training for each architecture. This distinction raises two important questions: First, how well do the condensed datasets generated by Gdem and Bonsai generalize across different architectures? Second, if we apply the condensation datasets produced by GCond and Gcsr for Gcn to other architectures, how significantly does it impact performance? We explore these questions in Table 5.
@@
Consistent with the earlier trend, Bonsai continues to outperform all baselines. We further observe that the performance gap between Bonsai and the baselines is wider in Gat and Gin compared to Gcn (Table 4). For GCond and Gcsr, this wider gap is not surprising since they are trained on the gradients of Gcn, which are expected to differ from those of Gin and Gat. Gin uses a Sum-Pool aggregation, while Gat employs attention to effectively dampen unimportant edges, with a potentially different eigen spectrum - the property that Gdem attempts to preserve. The approach of Bonsai, on the other hand, aligns more directly with the computational structure of Gnns. It focuses on analyzing the input space of computation trees and aims to represent as much of this space as possible within the given budget.
Table 5: Accuracies achieved by the various baselines on Gat and Gin.